### v1

# Nonparametric Estimation of Matching Efficiency and Elasticity on a Private On-the-Job Search Platform: Evidence from Japan, 2014-2024

Abstract
I use proprietary data from the online job scouting platform BizReach in Japan, spanning from 2014 to 2024, to estimate the matching function for high-skill employed workers on a private on-the-job search platform, employing a novel nonparametric approach developed by Lange and Papageorgiou (2020). This analysis is compared to the public off-the-job search platform, Hello Work. The results indicate that matching efficiency on the private platform is both more volatile and higher than that of the public platform, suggesting the increasing popularity of the private platform. The matching elasticity with respect to users consistently hovers around 0.75, while the elasticity with respect to vacancies reaches approximately 1.0, indicating a higher and more balanced elasticity compared to the Hello Work platform. Additionally, the study reveals evidence of industry-level heterogeneity on the private platform.
Keywords: matching efficiency, matching elasticities, on-the-job search, matching platform
JEL code: E24, J61, J62, J64
1 Introduction
On-the-job search plays a crucial role in labor reallocation, leading to wage and productivity improvements (Moscarini and Postel-Vinay 2017). In the U.S., job-to-job transitions account for one-third to one-half of all hires (Faberman et al. 2022). In contrast, Japan’s labor market has historically been characterized by long-term employment stability, with workers typically staying with a single employer for most of their careers. The proportion of employed individuals who changed jobs remains low (3.25 million), representing 4.8% of all employed workers in 2023. However, reflecting the Japanese government’s policies aimed at promoting job mobility, the number of employed individuals seeking to change jobs or explore new opportunities reached 10.35 million in 2023, representing 15.3% of the employed workforce. This marks the tenth consecutive period of growth and is the highest figure on record. [^1] According to the Labor Force Statistics Office of the Statistics Bureau, Ministry of Internal Affairs and Communications. Despite the growing importance of on-the-job search in understanding labor market dynamics in many countries—particularly through private platforms—empirical evidence on the extent, matching efficiency, and elasticity of this process remains limited. This contrasts sharply with the more abundant research on job search behavior among unemployed individuals.
This paper seeks to address this gap by providing new evidence on on-the-job search platforms for employed, high-skill workers. Using proprietary aggregate data from BizReach, a prominent private online job scouting platform in Japan, I estimate the matching function within the context of on-the-job search and recover matching efficiency and elasticity by applying a novel nonparametric approach developed by Lange and Papageorgiou (2020). The platform allows registered job seekers to upload resumes, become active, and receive scouting messages from companies and headhunters actively searching for specialized talent. Workers on this platform can apply to posted jobs or wait to be scouted, contrasting with conventional job search platforms where workers actively apply for vacancies. As of July 2024, more than 2.58 million employed and self-employed workers in Japan seeking to change jobs or explore new opportunities had registered on the platform, indicating a reasonable degree of representativeness of on-the-job job seekers. To compare this private platform with the public sector counterpart studied in Otani (2024), I incorporate data from the “Report on Employment Service” (Shokugyo Antei Gyomu Tokei), using month-level aggregate data to analyze the trends of matching unemployed workers with full-time vacancies via the public employment platform, Hello Work. This comparison offers insights into the differing features and outcomes between private and public job search platforms in Japan.
My results highlight significant differences in matching efficiency and elasticity between the Hello Work public employment platform and a private BizReach platform from 2014 to 2024. For Hello Work, matching efficiency remained relatively stable until 2021, after which it sharply declined. The elasticity with respect to unemployment was consistently lower, ranging from 0.1 to 0.4, while the elasticity with respect to vacancies gradually increased to 1.0, indicating that changes in vacancies had a greater impact on job matching than changes in unemployment. This offers different implications from the findings of Otani (2024), which used a broader time range. On the private platform, matching efficiency was highly volatile, peaking in 2016. Elasticities on the private platform were also more variable and generally higher than those observed in Hello Work. The elasticity with respect to users consistently hovered around 0.75, while the elasticity concerning vacancies steadily increased, reaching around 1.0 by 2024. This suggests a more responsive matching process to changes in the number of users and a more balanced elasticity between users and vacancies compared to Hello Work.
Industry-level analysis on the private platform reveals that the Consulting sector exhibits higher matching efficiency and responsiveness to labor market changes, especially after 2020, in contrast to the IT and Manufacturing sectors. The findings underscore industry heterogeneity, where sectors like IT and Internet are characterized by stable efficiency, similar to Manufacturing, while the Consulting sector experiences more variable matching efficiency. These sectoral insights provide a deeper understanding of labor market dynamics on private platforms, with clear contrasts to the public Hello Work system.
Overall, this paper offers quantitative insights from proprietary aggregate data into the matching function in on-the-job search labor markets, though it should be noted that the private platform analyzed may not fully represent the broader on-the-job search labor market in Japan.
1.1 Related literature
This paper contributes to three key areas of research: nonparametric matching functions, on-the-job search, and online job search platforms operated by private firms.
First, it adds to the empirical literature on the estimation of the matching function, a foundational component in macroeconomic models. Using a novel nonparametric approach developed by Lange and Papageorgiou (2020), I examine trends in matching efficiency in Japanese labor markets via an online job scouting platform. This method enables the identification and estimation of the matching function without imposing the standard independence assumption between matching efficiency and search efforts from either side of the labor market. This approach accommodates multiple types of job seekers. Lange and Papageorgiou (2020) highlight the positive correlation between efficiency and market structure variables like labor market tightness, which introduces a positive bias in vacancy elasticity estimates unless unobserved matching efficacy is accounted for. In traditional Cobb-Douglas matching function models, this unobserved factor is often ignored, resulting in potentially biased elasticity estimates. [^2] For example, Petrongolo and Pissarides (2001) summarize early aggregate studies using the Cobb-Douglas matching function, finding the match elasticity with respect to unemployment to be in the range of 0.5 to 0.7. In the context of Japan, Otani (2024) update the earlier findings of Kano and Ohta (2005), Kambayashi and Ueno (2006), and others, comparing results with international findings like Bernstein et al. (2022).
Second, this paper relates to the Japanese labor market literature, particularly in the context of the public off-the-job search platform, Hello Work. Otani (2024) estimate matching efficiency and mismatch in Japan’s Hello Work platform between 1972 and 2024, showing a declining trend in matching efficiency consistent with decreasing job and worker finding rates. The match elasticity with respect to unemployment is found to be between 0.5 and 0.9, while the elasticity concerning vacancies ranges from 0.1 to 0.4. In comparison, Otani and Kanayama (2024) apply a similar nonparametric method to estimate matching efficiency and elasticity in a privately-operated spot-worker platform, contrasting these findings with Hello Work’s part-time data. My paper complements and extends these findings by focusing on a high-skill, employed worker platform, filling an important gap in the literature. Table 1 provides an overview of recent empirical studies on Japanese labor markets between January 2014 and April 2024, contributing new evidence on matching efficiency, elasticity, and mismatch.
Table 1: Nonparametric estimation of the matching function in Japanese Labor markets
Paper Platform operator Worker status Vacancy type
Otani (2024) Administrative Unemployed Full-time, part-time
Otani and Kanayama (2024) Timee Unemployed Part-time, spot work
This paper BizReach Employed Full-time
The Japanese government’s efforts to enhance labor market flexibility have significantly contributed to the increase in on-the-job search activities. Initiatives such as the 2018 Work Style Reform Laws, aimed at reducing overwork, improving work-life balance, and promoting diverse career trajectories, have encouraged workers to explore new career opportunities without the societal stigma previously associated with job changes. These reforms have reshaped attitudes toward job mobility and created a more flexible labor market environment, enabling workers to transition between roles more easily while still employed (Yamamoto 2019, Owan 2017). This policy context underscores the relevance of quantitative studies on private on-the-job search platforms, such as those examined in this paper.
Second, this paper contributes to the literature on on-the-job search, a key aspect of labor search theory since the 1970s (Burdett 1978). Recent theoretical models, including those by Cahuc et al. (2006), Eeckhout and Lindenlaub (2019), and Bagger and Lentz (2019), emphasize the role of search effort in on-the-job search and its connection to job ladder dynamics. Empirical research, such as Mueller (2010) and Ahn and Shao (2017), has relied on data from the American Time Use Survey (ATUS) to document on-the-job search behaviors. However, due to the limitations of ATUS in capturing search outcomes, there is a gap in evaluating the efficiency of on-the-job search. Notably, Faberman et al. (2022) and Roussille and Scuderi (2023) provide crucial insights, with Faberman et al. (2022) focusing on the relationship between search effort and outcomes and Roussille and Scuderi (2023) exploring wage markdown using data from Hired.com. However, these studies lack long-term macro-level insights into matching function, efficiency, and elasticity. The proprietary data in this paper offers a unique advantage by allowing for the evaluation of matching efficiency in the on-the-job search labor market via a private platform.
Table 2: Literature on private online job search platforms (U.S)
Paper Period Platform Worker Vacancy
U.S.
Kuhn and Skuterud (2004) 98-00 Any U Both
Kuhn and Mansour (2014) 05-08 Any U Both
Faberman et al. (2022) 13-17 Any Both Both
Hershbein and Kahn (2018) 07,10–15 BGT U Both
Hershbein and Macaluso (2018) 07,10-17 BGT Both Both
Burke et al. (2020) 07,10-17 BGT Both Both
Adams-Prassl et al. (2024) 13-19 BGT Both Both
Schubert et al. (2024) 11-19 BGT Both Both
Bana (2021) 19-20 BGT Both Both
Greenwich.HR
Marinescu (2017) 07-11 CareerBuilder Both Both
Marinescu and Rathelot (2018) 12 CareerBuilder U Both
Marinescu and Wolthoff (2020) 12 CareerBuilder U Both
Azar et al. (2020) 16 CareerBuilder Both Both
Azar et al. (2022) 12 CareerBuilder U Both
Kroft and Pope (2014) 06 Craigslist Both Both
Sockin et al. (2021) 13-20 Glassdoor Both Both
Gadgil and Sockin (2020) 08-19 Glassdoor Both Full-time
Sockin (2022) 08-21 Glassdoor Both Both
Roussille and Scuderi (2023) NA Hired.com Both Both
Roussille (2024) NA Hired.com Both Full-time
Sinclair and Gimbel (2020) 14-19 Indeed Both Full-time
Maurya and Telang (2018) NA LinkedIn Both Both
Brenčič and McGee (2023) 06 Monster.com Both Both
Faberman and Kudlyak (2019) 10-11 SnagAJob Both, low-skill Part-time
Brown and Matsa (2016) 08-09 Private Both Both
Li et al. (2024) 16-19 Private Both, high-skill Full-time
Note: BGT=Burning Glass Technologies, U=Unemployed, E=Employed, Both=U&E
Table 3: Literature on private online job search platforms (Non US)
Paper Period Platform Worker Vacancy
Kuhn and Shen (2013) 08-10, China Zhaopin.com Both Both
Martins (2018) 12-16, Mozambique Emprego.co.mz Both Both
Banfi and Villena-Roldan (2019) 08-14, Chile trabajando.com Both Both
Banfi et al. (2019) 08-16, Chile trabajando.com Both Full-time
Adrjan and Lydon (2019) 14-19, Ireland Indeed Both Both
Kuhn et al. (2020) 10, China XMRC.com Both Full-time
Chen and Luo (2022) 18-19, China 51job.com Both Both
Arceo-Gomez et al. (2022) 18-20, Mexico OMBC Both Both
Choi et al. (2022) 10-20, Chile trabajando.com Both Both
Banfi et al. (2022) 10-19, Chile trabajando.com Both Full-time
Afridi et al. (2022) 19-21, India HelpersNearMe Both, blue Both
Kang and Shen (2022) 18-19, China XMRC.com Both Full-time
Barwick et al. (2023) 16-18, China Zhaopin.com Both Both
58.com
He et al. (2023) 23, China Private Both Both
Kiss et al. (2023) 22, South Africa SAYouth.mobi Both, low-skill Both
Brinatti et al. (2023) 19 & 20, global Private Both Both
Rouwendal and Koster (2023) 18, Netherlands Any Both Both
Grasso and Tatsiramos (2023) 17-19, Italy WollyBi Both Both
Escudero et al. (2024) 10-20, Uruguay BuscoJobs Both Both
Subramanian et al. (2024) 17-22, Pakistan Job Talash Both Both
Jensen (2024) 07-17, Denmark HBS Economics E Both
Benner1 et al. (2024) 17-22, Germany Private Both Both
Bassier et al. (2024) 17-19, UK Adzuna Both Both
Zvedelikova (2024) 18-19, Japan Doda Both Full-time
Otani and Kanayama (2024) 18-23, Japan Timee U, low-skill Part-time
This paper 14-23, Japan BizReach E, high-skill Full-time
Note: U=Unemployed, E=Employed, Both=U&E, OMBC=OOC Mundial, Bumeran, & CompuTrabajo
Third, this paper contributes to the expanding literature on online job search platforms. The analysis of job matching within real-world market institutions has gained prominence due to the increasing availability of data from online job platforms (Autor 2019), as summarized in Tables 2 and 3.333Examples include studies like Kuhn and Skuterud (2004), Kuhn and Mansour (2014), and Kroft and Pope (2014), which focus on worker status, while others, such as Kuhn and Shen (2013), Hershbein and Kahn (2018), Brown and Matsa (2016), and Azar et al. (2020), focus solely on vacancy data. Moreover, research like Banfi and Villena-Roldan (2019), Marinescu and Rathelot (2018), Marinescu and Wolthoff (2020), and Azar et al. (2022) incorporates both worker and vacancy information. Much of the literature emphasizes application-level or vacancy-level behavior to assess search behavior and wage elasticity. For instance, Faberman and Kudlyak (2019) leverage proprietary application-level data from an online job search engine to explore the relationship between search intensity and duration, primarily focusing on lower-skill, hourly jobs for employed and unemployed workers. Similarly, Kambayashi et al. (2023) estimate elasticities of application, interview attendance, and offer acceptance relative to posted wages using detailed process-level data from private job search and matching intermediary platforms in Japan, which became the most significant recruitment channel in 2023, as shown in Figure 1. In contrast, this paper adopts a broader macro-level perspective, evaluating the overall efficiency of the private matching platform. To the best of my knowledge, this is the first paper to estimate matching efficiency and elasticity in an online job scouting platform, offering relatively long-term insights into private online job search trends, complementing the micro-level studies.
Figure 1: Mid-Career Recruitment Channels Prioritized by Companies
Source: Cabinet Office, Government of Japan (2024). The author reproduces Figure 2 in “2024 Annual Economic and Fiscal Report” Chapter 2. The survey asks sampled firms whether they evaluate each channel or not.
2 Data
2.1 Data source
First, I use the Report on Employment Service (Shokugyo Antei Gyomu Tokei) for month-level aggregate data from January 2014 to April 2024 to examine trends in matching unemployed workers with vacancies via Japan’s public employment platform, Hello Work. These datasets include the number of job openings, job seekers, and successful job placements, primarily sourced from the Ministry of Health, Labour and Welfare (MHLW) of Japan, which regularly publishes monthly reports and statistical data on the Public Employment Security Office, commonly known as Hello Work. Hello Work plays a crucial role in Japan’s labor market by providing government-operated employment counseling, job placement services, and vocational training. It has been extensively used for estimating traditional Cobb-Douglas matching functions, as seen in studies like Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018), as well as nonparametric estimation Otani (2024). In this study, I focus on full-time workers to ensure consistency for comparison across different datasets. The chosen period provides a consistent timeframe for comparison with the following platform data.
Second, I utilize proprietary data from BizReach, a private job-scouting platform in Japan, to analyze trends in matching employed workers with vacancies. To maintain consistency with the Hello Work data, I include only “active” workers, defined as those who logged into the platform in a given month, excluding inactive registered users. Unlike Hello Work, BizReach caters to high-level professionals and executives, offering a premium job-scouting service. Candidates can either use the platform for free or pay a monthly subscription fee (approximately 40 U.S. dollars) to gain priority access to job opportunities and services.
The BizReach platform allows job seekers to upload resumes and receive scouting messages from companies or headhunters searching for specialized talent. This system encourages proactive recruitment, enabling direct communication between job seekers and employers. Users also gain insights into their market value through the scouts they receive, even if they are not actively searching for new opportunities. This platform’s focus on high-skill professionals contrasts with the broader services provided by Hello Work, which includes support for entry-level and part-time positions. While BizReach emphasizes efficiency in high-level recruitment, it may not be suitable for individuals seeking entry-level roles or more comprehensive career counseling services, which Hello Work provides.
2.2 Trend comparison
Figure 2: Trends of key variables: Hello Work full-time (left) vs platform (right) 2014-2024
Note: For confidentiality reasons, the y-axis levels in the right panels have been masked, making them not directly comparable with those in the left panels. Additionally, labor market tightness on the platform has not been reported to maintain confidentiality.
Figure 2 provides a comparative analysis of labor market dynamics between the Hello Work public employment platform (left panel) and a private scouting platform (right panel) from 2014 to 2024. For confidentiality reasons, the y-axis levels for the platform panels have been masked and are therefore not directly comparable with Hello Work panels.
In the Hello Work panel, the unemployment trend remains relatively stable, fluctuating around 1.5 million over the observed period, with a slight decline noted after 2020. The number of vacancies increases steadily, resulting in a moderate rise in labor market tightness, although this ratio stays below 1 throughout the period. Both the hiring count and the job-worker finding rates which are the numbers of matching hires divided by the number of vacancies and workers exhibit a downward trend, suggesting two possible explanations. The first is the existence of potential challenges, inefficiencies, or mismatches in job placements facilitated by the Hello Work platform. The second alternative explanation could be the growing presence of private job search platforms, which offer an alternative avenue for unemployed workers seeking career transitions. It is important to note, however, that the private platform analyzed in this paper primarily caters to employed workers, thus limiting direct evidence regarding its impact on unemployed workers.
The right panels in the updated figure illustrate labor market trends on the private platform for high-skilled employed workers from 2014 to 2024. There is a steady and consistent increase in employed users, especially starting around 2018. According to the Japanese Labor Force Survey, the total number of employed workers is 66 million in 2014 and 69 million in 2024. Additionally, the Labor Force Statistics Office reports that approximately 10.35 million employed individuals were seeking to change jobs or explore new opportunities in 2023. Thus, around 1% of the total employed workforce and nearly 5% of those actively exploring job changes have registered on the platform, indicating their engagement in on-the-job search activities. However, this does not necessarily imply active job pursuit for all registered users. Conversely, vacancies increase gradually but remain relatively modest, resulting in a low labor market tightness ratio, which is not reported above for confidentiality reasons. This low tightness ratio does not necessarily indicate a mismatch between job supply and demand on the platform, as registered workers may be passively exploring opportunities rather than actively searching for new positions.
The right panel of Figure (b) illustrates the hiring count, which shows a gradual and steady increase beginning around 2018. However, even with this upward trend, the overall hiring level masked in the figure remains relatively modest compared to Hello Work, particularly when considering the substantial rise in the number of users. This pattern may suggest that many users are engaging with the platform in a more passive manner rather than actively seeking new employment. Panel (c) presents the job-worker finding rates, with the worker-finding rate displaying slight fluctuations but maintaining a generally comparable level to that of Hello Work. Meanwhile, the job-finding rate remains consistently low throughout the period compared to the worker-finding rate. These patterns suggest a notable disparity between worker-side and vacancy-side matching probabilities, despite the marked growth in platform users and job postings.
3 Model
Our primary focus lies in analyzing matching efficiency and matching elasticity with respect to the number of registered workers and vacancies in the labor market, as facilitated by an online job scouting platform operated by a private firm in Japan. A matching function derived from search models plays a pivotal role in labor economics. [^4] See Pissarides (2000), Petrongolo and Pissarides (2001), and Rogerson et al. (2005) for further reference. The matching function operates on the premise of random search from both sides of the labor market, where job seekers represent labor supply and recruiters represent labor demand. Conceptually, this paper examines two independent matching functions on the off-the-job and on-the-job search labor markets. Examining interdependence between two matching functions theoretically is out of the scope of this paper.
To estimate the matching function and recover matching efficiency, I adopt the novel approach proposed by Lange and Papageorgiou (2020). [^5] Lange and Papageorgiou (2020) additionally incorporate search effort (Mukoyama et al. 2018) and a recruitment intensity index (Davis et al. 2013). The paper highlights two critical issues: the endogeneity of matching efficiency (Borowczyk-Martins et al. 2013) and the overly restrictive nature of the Cobb-Douglas specification, which assumes fixed matching elasticity. To address these limitations, Lange and Papageorgiou (2020) propose a nonparametric identification and estimation framework for matching efficiency under specific conditions that will be discussed later.
Let unscripted capital letters $(A,U,V)$ denote random variables, while time-specific realizations are subscripted by $t$. I consider the matching function $m_{t}(\cdot,\cdot)$, which maps period-$t$ users $U_{t}$, per-capita search efficiency/matching efficiency $A_{t}$, and vacancies $V_{t}$ into hires $H_{t}$. For the private platform analysis, $U_{t}$ represents the number of employed workers registered on the platform, whereas $U_{t}$ for Hello Work refers to unemployed workers registered in the public system. I assume a stationary data-generating process, with sufficient time-series data to treat the joint distribution $G:\mathbb{R}_{+}^{3}\rightarrow[0,1]$ of $\left(H_{t},U_{t},V_{t}\right)$ as observable. Additionally, I denote by $F(A,U)$ the joint distribution of $A$ and $U$.
I identify the matching function and the unobserved, time-varying matching efficiency, $A$. First, I assume that $V$ and $A$ are conditionally independent given $U$, i.e., $A\perp V\mid U$. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ exhibits constant returns to scale (CRS). [^6] To align with the original model of Matzkin (2003), the function $H=m(AU,V)$ can be reformulated as $H/U=m(A,V/U)$ under CRS, where $H/U$ and $V/U$ are the job-finding rate and market tightness, respectively. These two assumptions are commonly used in the literature. Applying the nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) demonstrates that the joint distribution $G(H,U,V)$ identifies $F(A,U)$ and the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point, denoted as $A_{0}$, within the support of $(A,U,V)$. [^7] In Otani (2024), I report finite sample performance and extend the methodology through Monte Carlo simulation. Simulation results with a sample size of $T=50$ indicate that the sample size in this paper is sufficient for accurately recovering matching efficiency. For practical issues, see Brancaccio et al. (2020b), which applies this approach to estimate a matching function in a trade model (Brancaccio et al. 2020a, 2023).
4 Estimation
Following Lange and Papageorgiou (2020), I begin by estimating $F(A_{0}|U)$ across the support of $U$. To achieve this, I use the distribution of hires conditional on users, $U$, and observed vacancies, $V$. Specifically, we have:
$F(A_{0}|\psi U_{0})=G_{H|U,V}(\psi H_{0}|\psi U_{0},\psi V_{0})\quad\text{for any arbitrary scalar }\psi,$
$F(\psi A_{0}|\lambda U_{0})=G_{H|U,V}(\psi H_{0}|\lambda U_{0},\psi V_{0})\quad\text{where }\lambda>0\text{ is a scaling factor},$
where $F(A_{0}|\psi U_{0})$ and $G_{H|U,V}$ represent the respective conditional distributions. By varying the parameters $(\psi,\lambda)$, I can trace out $F(A|U)$ across the entire support of $(A,U)$.
Given that my data is finite, I rely on an estimate of $G_{H|U,V}$ for the constructive estimator. Consider an arbitrary point $(H_{\tau},U_{\tau},V_{\tau})$. To obtain $G(H_{\tau}|U_{\tau},V_{\tau})$, I calculate the proportion of observations with fewer hires than $H_{\tau}$, taken from observations proximate to $(U_{\tau},V_{\tau})$ in the $(U,V)$-space. In practice, this is done by averaging across all observations, assigning smaller weights to those with values $(U_{t},V_{t})$ distant from $(U_{\tau},V_{\tau})$ via a kernel that discounts distant observations. The resulting estimate of $F(\psi A_{0}|\lambda U_{0})=G_{H|U,V}(\psi H_{0}|\lambda U_{0},\psi V_{0})$ is expressed as:
$\hat{F}(\psi A_{0}|\lambda U_{0})=\sum 1(H_{t}<\psi H_{0})\kappa(U_{t},V_{t},\lambda U_{0},\psi V_{0}),$
where $\kappa(.)$ denotes a bivariate normal kernel with bandwidth 0.01.
Once the distribution function $F(A|U)$ is recovered, I invert $F(A_{t}|U_{t})$ to derive $A_{t}$ for all observations in the dataset, using:
$A_{t}=F^{-1}(G(H_{t}|U_{t},V_{t})|U_{t}),$
Finally, I recover the matching function as:
$m(A_{t},U_{t})=G^{-1}(F(A_{t}|U_{t})|U_{t}).$
To compute matching elasticities, I employ a LASSO regression, projecting hires onto the original and squared values of vacancies and users, interacted with implied matching efficiency. The resulting estimates approximate the derivatives of the matching function with respect to vacancies and users, interacted with implied matching efficiency. This provides an estimate of the elasticity of the matching function. [^8] The matching elasticity with respect to users $\frac{d\log m(AU,V)}{d\log U}=\frac{dm(AU,V)}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{dAU}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{AU}{H}=\frac{d\log m(AU,V)}{d\log AU}$ is obtained from the regression coefficient of $H$ on $AU$ and multiplying it by $\frac{AU}{H}$. Concretely, I approximate $m$ by the second order polynomial $AU$ and $V$.
5 Results
5.1 Matching efficiency and elasticity in the platform
Figure 3: Hello Work full-time vs platform 2014-2024
Figure 3 presents a comparison of matching efficiency and matching elasticity between the Hello Work public employment platform (left panel) and a private scouting platform (right panel) from 2014 to 2024. In the Hello Work platform, as depicted in panel (a), matching efficiency remains relatively stable around the baseline (normalized to January 2014) until 2021 but sharply declines thereafter. Panel (b) illustrates the matching elasticities with respect to unemployment and vacancies. The elasticity with respect to unemployment remains consistently lower, around 0.4, while the elasticity with respect to vacancies shows a gradual increase, reaching values near 1.0 by 2022. This suggests that changes in vacancies have a larger impact on the number of matches than changes in unemployment. [^9] The estimated elasticities differ from those reported in Otani (2024), largely due to the different time horizons considered. Specifically, Otani (2024) includes data from 1972 to 2024, a period that encompasses various economic booms and busts, which likely captures a broader range of labor market dynamics and affects the elasticity estimates.
On the private scouting platform (right panel), panel (a) shows significantly higher volatility in matching efficiency, especially after 2016, where it peaks sharply above 900 before declining and stabilizing around 400 by 2024. This indicates that the private platform experienced large fluctuations in its ability to efficiently match job seekers with vacancies, potentially due to shifts in demand or changes in the platform’s user base. Panel (b) displays the matching elasticities with respect to users and vacancies on the private platform. The elasticity with respect to users remains consistently around 0.75, while the elasticity with respect to vacancies is generally higher, rising steadily and reaching around 1.0 by 2024. This suggests a more responsive matching process to changes in the number of vacancies. The higher volatility and responsiveness of the private platform, along with a more balanced elasticity between users and vacancies, highlight a key distinction from Hello Work, where the dynamics are comparatively more stable but less responsive.
5.2 Industry-level matching efficiency and elasticity in the platform
Figure 4: Industry-level results on Platform 2014-2024
Note: For confidentiality reasons, the y-axis levels are masked.
Figure 4 illustrates labor market dynamics across three sectors—Consulting, IT and Internet, and Manufacturing—on a private job platform from 2014 to 2024. These categories represent high-skill full-time employed workers, though they differ from those used by Hello Work. For confidentiality reasons, the precise numbers of users, vacancies, hires, and the y-axis levels are not disclosed, but the trends provide valuable insights into sector-specific patterns.
Panel (a) shows that labor market tightness ($V/U$) has generally increased across all three sectors from 2014 to 2024. Consulting and IT & Internet sectors exhibit higher and more volatile tightness ratios, particularly after 2020, reflecting more job openings per user. In contrast, the Manufacturing sector shows a lower and more stable tightness ratio, indicating fewer vacancies relative to job seekers in this field.
Panel (b) depicts matching efficiency, with the IT and Internet sector exhibiting the highest variability and levels, especially after 2020. Panels (c) and (d) illustrate the matching elasticities with respect to users and vacancies, respectively. The Consulting sector shows negative elasticity with respect to users early on, likely due to a small user base, but this increases sharply over time. Meanwhile, elasticity with respect to vacancies slightly decreases. IT and Internet and Manufacturing sectors show more stable elasticity patterns. The overall trends highlight significant industry-level heterogeneity, with Consulting experiencing the most dynamic growth, while IT and Internet and Manufacturing follow more stable but slower trends.
6 Conclusion
This paper uses proprietary data from BizReach, an online job scouting platform in Japan, spanning from 2014 to 2024, to estimate the matching function for high-skill employed workers in a private on-the-job search platform. The results are compared to a public off-the-job search platform, specifically targeting unemployed workers seeking full-time jobs. Findings suggest that matching efficiency on the private platform is more volatile but generally higher than on the public platform, highlighting the increasing reliance on private platforms for high-skill job searches. Furthermore, the private platform demonstrates a higher matching elasticity with respect to unemployment (between 0.6 and 0.8) compared to the public platform, while the elasticity with respect to vacancies is similar (between 0.8 and 1.1). This indicates a more balanced responsiveness to changes in both users and vacancies on the private platform, compared to the more stable but less dynamic public platform, Hello Work.
Additionally, industry-level heterogeneity is evident across both platforms, reflecting differing labor market dynamics by sector. However, while this paper provides key insights into the matching function for on-the-job searches, the analysis may not fully represent the broader on-the-job search labor market, particularly for non-high-skill workers in Japan. Moreover, the standard assumption of homogeneity among workers and vacancies may overlook important nuances. Future research should focus on expanding this analysis to other private platforms and exploring individual-level behavior, as discussed in studies like Kambayashi et al. (2023) and Roussille and Scuderi (2023), to provide a more comprehensive understanding of labor market dynamics.